# Condensates in aggregation processes with input: A new power law and intermittency

### v2

@@ VI Intermittency
Finally, we point out that in the steady state of Takayasu model with open boundary conditions, the exits of condensates with mass $\sim O(L^{3})$, lead to large crashes of the total mass $M(t)$ of the system in time (see Fig. 6). The crashes are followed by a build-up of $M(t)$ due to input, followed again by crashes, and so on in an intermittent fashion, reminiscent of dragon king events Sornette and Ouillon (2012); Mikaberidze et al. (2023).
@@ VII Conclusion
In a broader vein, it would be interesting to look for strong imprints of condensates on the static and dynamic properties of other systems obeying aggregation-diffusion dynamics. On the theoretical end, the conserved mass aggregation model Majumdar et al. (1998, 2000a) and the in-out model Majumdar et al. (2000b); Rajesh (2004) both exhibit condensates, and are therefore expected to share some of the signatures discussed above. Experimentally, biomolecular aggregates which are described within the Takayasu framework of aggregation and input Brangwynne et al. (2011); Berry et al. (2018); Lee et al. (2023), would be promising candidates to look for these signatures.
@@ VIII Acknowldegement
We thank Arghya Das and Himani Sachdeva for helpful discussions, and Deepak Dhar for a critical reading of the manuscript. We also acknowledge Eli Barkai and Punyabrata Pradhan for useful comments. M. B. acknowledges the support of the Indian National Science Academy (INSA). We acknowledge the support of the Department of Atomic Energy, Government of India, under Project Identification No. RTI4007.

### v3

@@ VI Intermittency
Finally, we point out that in the steady state of Takayasu model with open boundary conditions, the exits of condensates with mass $\sim O(L^{3})$, lead to large crashes of the total mass $M(t)$ of the system in time (see Fig. 6). The crashes are followed by a build-up of $M(t)$ due to input, followed again by crashes, and so on in an intermittent fashion, reminiscent of dragon king events [25, 26].
@@ VII Conclusion
In a broader vein, it would be interesting to look for strong imprints of condensates on the static and dynamic properties of other systems obeying aggregation-diffusion dynamics. On the theoretical end, the conserved mass aggregation model [18, 19] and the in-out model [31, 32] both exhibit condensates, and are therefore expected to share some of the signatures discussed above. Experimentally, biomolecular aggregates that are described within the Takayasu framework of aggregation and input [4, 5, 6], would be promising candidates to look for these signatures.
@@ VIII Acknowldegement
We thank Arghya Das and Himani Sachdeva for helpful discussions, and Deepak Dhar for a critical reading of the manuscript. We also acknowledge Eli Barkai and Punyabrata Pradhan for their useful comments. M.B. acknowledges the support of the Indian National Science Academy (INSA). We acknowledge the support of the Department of Atomic Energy, Government of India, under Project Identification No. RTI4007.